# ROTATE: Regret-driven Open-ended Training for Ad Hoc Teamwork

### v1

@@ 5 An Open-Ended Learning Perspective on Ad Hoc Teamwork
$\text{CR}(\psi^{\text{eval}},\Pi^{\text{eval}},\pi^{\text{ego}})=\mathbb{E}_{\pi^{-i}\sim{\psi^{\text{eval}}(\Pi^{\text{eval}})},s_{0}\sim p_{0}}\left[\text{CR}((\pi^{\text{ego}},\pi^{-i},s_{0})\right].$ (6)
@@
where we re-emphasize that $\Pi^{-i}$ denotes the set of all competent and non-adversarial [50] joint teammate policies. Limiting the considered joint policies is important since teams that always perform poorly against any good-faith $\pi^{\text{ego}}$ are unlikely to be encountered in coordination scenarios and may induce unnecessary learning challenges for RL-based AHT learning algorithms.
Finding $\pi^{\text{ego}}$ that achieves zero worst-case regret is equivalent to finding an ego agent that achieves the best-response return with any joint teammate policy $\pi^{-i}$. If such a $\pi^{\text{ego}}$ exists, then this AHT agent would maximize Eq. 2 for any $\psi^{\text{eval}}$ and $\Pi^{\text{eval}}$—however, existence is not guaranteed [30]. In practice, we are content with minimizing the worst-case regret. While minimizing worst-case regret still applies to AHT problems with more than one teammate at a time, note that we limit our method for optimizing Eq. 7 and our experiments to two-player, fully observable AHT games.
@@
6 ROTATE: Regret-driven Open-ended Training for Ad hoc TEamwork
This section presents our regret-driven, open-ended AHT algorithm, ROTATE. We first describe the teammate generation procedure in Section 6.1, particularly focusing on motivating the objective we used to generate teammate policies. Next, we provide details of the ego agent update method in Section 6.2. The Appendix provides ROTATE’s pseudocode and additional implementation details.
@@ 6.1 ROTATE Teammate Generator
The teammate generator produces teammate policies that maximize the regret of $\pi^{\text{ego}}$. By updating $\pi^{\text{ego}}$ to minimize its regret against the regret-maximizing teammate policy, we aim to decrease the worst-case cooperative regret of $\pi^{\text{ego}}$ (Eq. 7). Since measuring cooperative regret requires estimating the performance of a generated $\pi^{-i}$ when collaborating with $\text{BR}(\pi^{-i})$, we jointly train policies for $\pi^{-i}$ and an approximation of $\text{BR}(\pi^{-i})$ using the Proximal Policy Optimization (PPO) algorithm [48].
Before detailing our alternative objectives for training $\pi^{-i}$, we first introduce the different interactions that provide the experience to train $\pi^{-i}$. Let self-play (SP) refer to teammate and best response interactions, cross-play (XP) refer to teammate and ego agent interactions, and cross-play continued by self-play (SXP) refer to an interaction where the teammate is first interacting with the ego agent, but switches at a random timestep $t$ to interacting with the best response. We train $\pi^{-i}$ based on states sampled from SP, XP, and SXP. Let $d(\pi^{1},\pi^{2};p_{0})$ denote the state visitation distribution when $\pi^{1}$ and $\pi^{2}$ interact based on a starting state distribution $p_{0}$. To denote the state visitation distributions for these interactions, we use the following shorthand:
@@
This section considers two teammate policy generation objectives that differ in the data source used to optimize the objective, and are illustrated in Fig. 2.
Figure 2: Per-trajectory regret vs per-state regret.
The first objective generates $\pi^{-i}$ based on maximizing per-trajectory regret of $\pi^{\text{ego}}$, which only maximizes regret of trajectories starting from the initial state distribution:
From the definition of CR (Eq. 5), optimizing per-trajectory regret amounts to maximizing the expected returns of SP interactions from the initial state distribution $p_{0}$, $\mathbb{E}_{s\sim p_{0}}\left[V(s|\pi^{-i},\text{BR}(\pi^{-i}))\right]$, and minimizing the expected returns of XP interactions from $p_{0}$, $\mathbb{E}_{s\sim p_{0}}\left[V(s|\pi^{-i},\pi^{\text{ego}})\right]$. This objective resembles the adversarial diversity metric optimized in prior teammate generation work [9, 41, 14].
While the per-trajectory regret is the same as the regret objective optimized in Eq. 7, optimizing it naively leads to generating teammates with undesirable self-sabotage behaviors. To minimize the expected returns from cross-play interaction, self-sabotaging policies choose actions leading to low returns in states outside the support of $p_{\text{SP}}$ [14], including the states visited during the collaboration between $\pi^{-i}$ and $\pi^{\text{ego}}$, $s\sim p_{\text{XP}}$. The lack of high reward signals makes it challenging for the ego agent to learn to collaborate with $\pi^{-i}$ using reinforcement learning.
The second objective discourages the emergence of self-sabotaging policies by optimizing a per-state regret objective defined as:
$\underset{\pi^{-i}}{\max\ }(\mathbb{E}_{s\sim p_{\text{XP}}}[\text{CR}(\pi^{\text{ego}},\pi^{-i},s)]+\mathbb{E}_{s\sim p_{\text{SXP}}}[V(s|\pi^{-i},\text{BR}(\pi^{-i}))]+\mathbb{E}_{s\sim p_{\text{SP}}}[V(s|\pi^{-i},\text{BR}(\pi^{-i}))]).$ (10)
The first term in Expr. 10 encourages discovering $\pi^{-i}$ for which the ego agent policy has a high room for improvement. Meanwhile, the second term trains $\pi^{-i}$ to act as if it interacts with its best response policy, even from starting states encountered during XP interactions. This encourages $\pi^{-i}$ to act in good faith by having at least a partner policy that can collaborate well with $\pi^{-i}$ when starting from states in $p_{\text{XP}}$. Finally, the last term encourages $\pi^{-i}$ to cooperate with its best response. This enables consistently generating competent teammates during open-ended learning, which is essential as stated in Section 5.
While obtaining states from $p_{\text{SP}}$ and $p_{\text{XP}}$ is straightforward, states from $p_{\text{SXP}}$ can be tricky to collect depending on the implementation of an AHT environment. If an environment supports resetting to any arbitrary state, then states encountered during XP interaction can be stored and used as the initial state for SP interactions. Otherwise, we can use a data collection strategy that first samples a random timestep $t$, runs XP interaction until timestep $t$, and finally switches to SP interaction afterwards [46]. Only data gathered after timestep $t$ should be used to compute objectives based on $p_{\text{SXP}}$.
@@ 6.2 ROTATE Ego Agent Update
[…]
@@ 7 Experimental Results
This section presents the empirical evaluation of ROTATE compared to baseline methods, across six cooperative tasks. The main research questions are:
@@
  • RQ2: Does per-state regret improve over trajectory level regret and mixed-play regret? (Yes)
@@
  • RQ4: Is the population generated by ROTATE useful for training an independent ego agent? (Yes)
Supplemental results and analysis are provided in the Appendix. The code for the experiments in this paper is available at https://github.com/carolinewang01/rotate.
@@ 7.1 Experimental Setup
ROTATE is evaluated on six tasks: Level-Based Foraging (LBF) [2], and the five classic layouts from the Overcooked suite [8]: Cramped Room (CR), Asymmetric Advantages (AA), Counter Circuit (CC), Coordination Ring (CoR), and Forced Coordination (FC). All six tasks are fully cooperative benchmark tasks with a variety of possible coordination conventions, and are commonly used within the AHT literature [2, 12, 38]. In our LBF configuration, two agents must navigate to three apples that are randomly placed within a gridworld, and cooperate to pick up the apples. In all Overcooked tasks, two agents collaborate in a gridworld kitchen to cook and deliver onion soup, where the main difference between tasks is the kitchen layout. All experiments were implemented with JAX  [7], so we use JAX re-implementations of the LBF and Overcooked tasks [5, 45].
@@
As our method is most closely related to methods from unsupervised environment design (UED) and teammate generation, we compare against two UED methods adapted for AHT (PAIRED [15], Minimax Return [36, 54]) and three teammate generation methods (Fictitious Co-Play [51], BRDiv [41], CoMeDi [46]). While curator-based methods such as PLR [24, 25] are popular in UED, we do not compare against them as they are orthogonal to ROTATE [17, 54, 11]. Similarly, we do not compare against AHT algorithms that propose techniques to improve ego learning [3]. Each baseline is described in App. B, along with implementation details. For fair comparison, all open-ended and UED methods were trained for a similar number of environment interactions, or until best performance on the evaluation set. All teammate generation approaches were ran using a similar number of environment interactions as their original implementations, as scaling them up to use a similar number of steps as the open-ended approaches proved challenging (see discussion in App. B). All results are reported with three trials.
@@
7.2 Results
This section presents empirical analysis addressing the four research questions introduced at the beginning of Section 7. Supplemental analysis considering alternative regret-based objectives, breaking down performance by evaluation teammate type, and learning curves for all variants of ROTATE are provided in App. C.
To evaluate the generalization capabilities of ROTATE, we compare its performance against baselines on $\Pi^{\text{eval}}$. Fig. 3(a) compares the normalized mean returns for ROTATE and all baseline methods across the six tasks. The results show that ROTATE significantly outperforms all baselines on 5/6 tasks.
We also observe that Minimax Return performs surprisingly well in AA, which we hypothesize is due to AA’s particular characteristics. In AA, agents operate in separated kitchen halves, possessing all necessary resources for individual task completion, with shared access to pots on the dividing counter being the only shared resource. Consequently, a fully adversarial partner has limited methods to sabotage the ego agent. [^2] Agent teams may still achieve higher returns through effective coordination on AA, due to layout asymmetry. In the “left” kitchen, the delivery zone is adjacent to the pots while the onions are farther, while in the “right” kitchen, the opposite is true. Thus, an optimal team consists of the “left” agent delivering finished soup, and the “right” agent placing onions in the pots. However, on LBF and FC, where coordination is crucial to obtain positive return on the tasks, Minimax Return is the worst performing baseline.
@@ 7.2 Results
RQ2: Does per-state regret improve over trajectory regret? (Yes)
As discussed in Section 6, we propose that teammates should maximize per-state regret rather than per-trajectory regret to mitigate the emergence of self-sabotage behaviors. Here, we compare ROTATE where the teammate maximizes per-state regret (ours) to ROTATE where the teammate maximizes per-trajectory regret. All configurations other than the teammate’s policy objective are kept identical, including the data used to train the teammate value functions. Fig. 3(b) shows that ROTATE with per-state regret outperforms ROTATE with per-trajectory regret on all tasks except AA, confirming the superiority of per-state regret. As discussed in RQ1, we observe that AA is a layout where an ego agent is less susceptible to sabotage, due to the separated kitchen layout. App. C.2 presents additional experiments testing ROTATE with CoMeDi-style mixed-play rollouts, and alternative methods to compute per-state regret—ultimately finding that ROTATE outperforms all variations.
@@
RQ4: Is the population generated by ROTATE useful for training an independent ego agent? (Yes)
Two-stage AHT algorithms first generate a population of teammates, and next train an ego agent against the population. Although ROTATE’s teammate generation mechanism relies on the learning process of a particular ego agent, here, we investigate whether the population generated by ROTATE is useful for training independently generated ego agents. Fig. 5(b) (presented in the Appendix) compares the mean evaluation returns of the ROTATE ego agent against the mean evaluation returns of an independently trained ego agent that was trained using the same configuration as ROTATE. In 3/6 tasks, the ROTATE ego agent outperforms the independently trained ego agent, while in two tasks, the two ego agents perform similarly (LBF and FC). Thus, the experiment suggests that the ROTATE population is a useful population of teammates even independent of the particular ego agent generated. The strong performance of the independently trained ego agent is unsurprising given that it has two advantages over the ROTATE ego agent. First, the independently trained ego agent faces a stationary distribution of training teammates compared to ROTATE, which faces a nonstationary distribution caused by the population growing over learning iterations. Second, the independently trained ego agent interacts with all teammates uniformly throughout training, while the ROTATE ego agent only trains against earlier teammates for more iterations than later teammates.

### v2

@@ 5 An Open-Ended Learning Perspective on Ad Hoc Teamwork
$\text{CR}(\psi^{\text{eval}},\Pi^{\text{eval}},\pi^{\text{ego}})=\mathbb{E}_{\pi^{-i}\sim{\psi^{\text{eval}}(\Pi^{\text{eval}})},s_{0}\sim p_{0}}\left[\text{CR}(\pi^{\text{ego}},\pi^{-i},s_{0})\right].$ (6)
@@
where we re-emphasize that $\Pi^{-i}$ denotes the set of all competent and non-adversarial [50] joint teammate policies. Limiting the considered joint policies is important, as teams that consistently perform poorly against any $\pi^{\text{ego}}$ are unlikely to be encountered in coordination scenarios and may introduce unnecessary learning challenges for RL-based AHT learning algorithms.
Finding $\pi^{\text{ego}}$ that achieves zero worst-case regret is equivalent to finding an ego agent that achieves the best-response return with any joint teammate policy $\pi^{-i}$. If such a $\pi^{\text{ego}}$ exists, then this AHT agent would maximize Eq. 2 for any $\psi^{\text{eval}}$ and $\Pi^{\text{eval}}$. However, its existence is not guaranteed [30]. In practice, we are content with minimizing the worst-case regret. While minimizing worst-case regret still applies to AHT problems with more than one teammate, we limit our method for optimizing Eq. 7 and our experiments to two-player, fully observable AHT games.
@@
6 Regret-driven Open-ended Training for Ad Hoc Teamwork
This section presents our regret-driven, open-ended AHT algorithm, ROTATE. We first describe the teammate generation procedure in Section 6.1, focusing on motivating the objective used to generate teammate policies. Next, we describe the ego agent update method in Section 6.2. App. A provides ROTATE’s pseudocode and a more detailed implementation description.
@@ 6.1 ROTATE Teammate Generator
Given a fixed $\pi^{\text{ego}}$, ROTATE’s teammate generator seeks to discover a teammate policy that maximizes the cooperative regret of $\pi^{\text{ego}}$. Maximizing cooperative regret requires estimating the teammate policy, $\pi^{-i}$, and its best response policy, $\text{BR}(\pi^{-i})$. In the following, we abbreviate $\text{BR}(\pi^{-i})$ to BR for brevity. ROTATE’s teammate generator estimates both policies using the Proximal Policy Optimization (PPO) algorithm [48].
The per-trajectory regret of $\pi^{\text{ego}}$ (i.e., the inner objective of Eq. 7) is the regret from trajectories starting from the initial state distribution:
Eq. 8 resembles the objectives used in past UED [56, 17] and the teammate generation literature [42, 10] to generate tasks or teammate policies. Recent work demonstrates that maximizing per-trajectory regret is prone to yielding self-sabotaging teammates [16, 46]. Maximizing the cooperative regret only from $s_{0}\sim p_{0}$ implicitly encourages $\text{BR}(\pi^{-i})$ to select actions leading to future states that are distinguishable from those encountered during the interaction between $\pi^{-i}$ and $\pi^{\text{ego}}$. When encountering future states from interactions with $\pi^{\text{ego}}$, $\pi^{-i}$ ends up choosing actions that sabotage cooperation by minimizing the teammate’s returns against $\pi^{-i}$. Thus, training $\pi^{\text{ego}}$ to minimize regret (i.e., by maximizing the expected returns) when collaborating with $\pi^{-i}$ using RL becomes challenging because $\pi^{-i}$ actively chooses actions that undermine collaboration.
We mitigate the emergence of self-sabotage by training $\pi^{-i}$ to maximize two objectives across states sampled from different state visitation distributions. These state visitation distributions result from: (i) teammate and BR interactions (self-play, SP), (ii) teammate and ego agent interactions (cross-play, XP), and (iii) cross-play continued by self-play interactions (SXP), where the teammate is first interacting with the ego agent, but switches at a random timestep $t$ to interacting with its BR. Let $d(\pi^{1},\pi^{2};p)$ denote the state visitation distribution when $\pi^{1}$ and $\pi^{2}$ interact based on a starting state distribution $p$. We use the following shorthand to denote the SP, XP, and SXP state visitation distributions:
@@
Based on these distributions, we define the following per-state regret objective for training $\pi^{-i}$:
$\underset{\pi^{-i}}{\max\ }\left(\mathbb{E}_{s\sim 0.5\text{ }p_{\text{XP}}+0.5\text{ }p_{\text{SP}}}\left[\text{CR}(\pi^{\text{ego}},\pi^{-i},s)\right]+\mathbb{E}_{s\sim p_{\text{SXP}}}\left[V(s|\pi^{-i},\text{BR}(\pi^{-i}))\right]\right).$ (10)
The difference between the per-trajectory and per-step regret objectives is visualized in Figure 2. Both terms in the per-state regret objective discourage adversarial behavior from $\pi^{-i}$. The first term in Expr. 10 corresponds to the ego agent’s regret starting from both SP and XP states. Estimating regret from XP and SP requires collecting SXP data as well as an analogous type of data called XSP (SP continued by XP interactions), as detailed in App. A.2. In general, optimizing the ego agent’s regret encourages discovering $\pi^{-i}$ for which the ego agent policy has a high room for improvement. Optimizing regret starting from XP states requires $\pi^{-i}$ to be able to coordinate with its BR starting from any state encountered during interactions with the ego agent, thus preventing $\pi^{-i}$ from irrecoverably sabotaging an interaction. On the other hand, optimizing regret from SP states requires $\pi^{-i}$ to be able to decrease the return of the ego agent starting from any state encountered during interactions between the teammate and the BR, thus disincentivizing the emergence of unconditional cooperation signals. Finally, we find that training $\pi^{-i}$ to collaborate well with its BR even during SXP interactions helps ensure that $\pi^{-i}$ is a good-faith collaborator with at least one partner.
Figure 2: Teammate policy optimization objectives: per-trajectory regret vs per-state regret.
While obtaining states from $p_{\text{SP}}$ and $p_{\text{XP}}$ is straightforward, states from $p_{\text{SXP}}$ and $p_{\text{XSP}}$ are collected using either environment resetting or policy switching. Using SXP as an example, if an environment supports resetting to any arbitrary state, then states encountered during XP interaction can be stored and used as the initial state for SP interactions. Otherwise, we may sample a random timestep $t$, run XP interaction until timestep $t$, and then switch to SP interaction [46]. Only data gathered after timestep $t$ should be used to compute objectives based on $p_{\text{SXP}}$.
@@ 6.2 ROTATE Ego Agent Update
At each iteration, ROTATE creates a teammate that attempts to discover cooperative weaknesses of the previous iteration’s ego agent, by maximizing its per-state regret. To allow the ROTATE ego agent to improve its robustness over time and reduce the possibility that it forgets how to cope with previously generated teammates, the ROTATE ego agent maintains a population buffer of generated teammates. During the ego agent update phase of each iteration, the ROTATE ego agent is trained using PPO [48] against teammates sampled uniformly from the population buffer. We find experimentally that for the ego agent to learn effectively against the nonstationary population buffer, it is important to define a lower entropy coefficient and learning rate than when training the teammate and BR agents (typically in the range of $1\text{\times}{10}^{-4}$ for the entropy coefficient and $1\text{\times}{10}^{-5}$ for the learning rate).
@@ 7 Experimental Results
This section presents the empirical evaluation of ROTATE compared to baseline methods, across one illustrative matrix game and six benchmark tasks. Supplemental results, implementation details, and code link are provided in the Appendix. The main research questions are:
@@
  • RQ2: Does per-state regret mitigate sabotage and improve generalization to unseen teammates compared to per-trajectory regret? (Yes)
@@ 7.1 Experimental Setup
ROTATE is evaluated on a didactic matrix game and six benchmark tasks. For clarity, the matrix game is described with the corresponding results. The benchmark tasks are Level-Based Foraging (LBF) [2], and the five classic layouts from the Overcooked suite [9]: Cramped Room (CR), Asymmetric Advantages (AA), Counter Circuit (CC), Coordination Ring (CoR), and Forced Coordination (FC). All six tasks are cooperative, permit a variety of possible conventions, and are commonly used within the AHT literature [2, 13, 39]. In LBF, two agents must navigate to apples that are randomly placed within a gridworld, and cooperate to pick up the apples. In all Overcooked tasks, two agents collaborate in varying gridworld kitchen layouts to prepare dishes. All experiments were implemented with JAX  [8].
@@
As our method is most closely related to methods from UED and teammate generation, we compare against two UED methods adapted for AHT (PAIRED [17], Minimax Return [36, 54]) and three teammate generation methods (Fictitious Co-Play [51], BRDiv [41], CoMeDi [46]). While curator-based methods such as PLR [24, 25] are prevalent in UED, we do not compare against them as they are orthogonal to ROTATE [19, 54, 12]. Similarly, we do not compare against AHT algorithms for ego learning [3]. Each baseline is described in detail in App. B. For fair comparison, all open-ended and UED methods were trained for a similar number of environment interactions, or until best performance on the evaluation set. All teammate generation approaches were ran using a similar number of environment interactions as their original implementations, as scaling them up to use a similar number of steps as the open-ended approaches proved challenging (see discussion in App. B). All results are reported with three trials.
@@
7.2 Results
This section addresses the three research questions introduced at the beginning of Section 7. Supplemental analysis considering alternative regret-based objectives, independent utility of the ROTATE population, performance breakdowns by evaluation teammate, learning curves for all variants of ROTATE, and a human proxy evaluation on Overcooked, are provided in App. C.
We evaluate ROTATE’s generalization capabilities by comparing its performance against baselines on $\Pi^{\text{eval}}$. Fig. 3(a) compares the normalized mean returns for ROTATE and baseline methods across the six tasks. The results show that ROTATE significantly outperforms all baselines on 5/6 tasks.
Minimax Return performs surprisingly well in AA, which may be attributed to AA’s particular characteristics. In AA, agents operate in separated kitchen halves, possessing all necessary resources for individual task completion, with pots on the dividing counter being the only shared resource. A team where both agents act fully independently may achieve high returns–albeit coordination leads to still higher returns. [^2] Optimal behavior in AA still requires effective coordination due to layout asymmetry. In the “left” kitchen, the delivery zone is adjacent to the pots while the onions are farther, while in the “right” kitchen, the opposite is true. Thus, an optimal team consists of the “left” agent delivering finished soup, and the “right” agent placing onions in the pots—and indeed, we observe that teams of IPPO agents converge to this behavior. Visualizing the trained policies reveals that the adversarial teammate trained by Minimax Return cannot drive the ego agent’s return to zero, and does not prevent the ego agent from learning how to perform the task independently. However, on LBF and FC, where coordination is crucial to obtain positive returns, Minimax Return is the worst-performing baseline.
RQ2a: Does per-state regret mitigate sabotage compared to per-trajectory regret? (Yes)
H T S
H 1 0 -1
T 0 1 -1
S -1 -1 -1
Table 1: Payoff matrix for the sabotage game.
We design a simple sabotage game to investigate the whether the per-state regret objective leads to teammate policies that sabotage less often compared to the per-trajectory regret objective. The sabotage game is a fully cooperative, iterated matrix game with payoff matrix shown in Table 1. Each agent observes a game state that consists of the complete history of joint actions. Both agents have three actions: H, T, and S(abotage). The first two actions lead to two possible cooperative outcomes, while the last action leads to a reward of 1-1 if selected by either agent and immediate episode termination. Thus, the last action corresponds to sabotaging the team’s payoffs. By default, the game lasts for five timesteps.
We train ROTATE with both per-state and per-trajectory regret. To measure the extent to which the learned teammate policies engage in sabotage, we enumerate the 341 non-terminal states in the game and measure the probability of the sabotage action at each state for the last generated teammate policy. Fig. 4 shows that ROTATE with per-state regret has a near-zero probability of taking the sabotage action at all non-terminal states, while the per-trajectory regret objective leads to over a third of states that have P(S) near $1.0$.
Figure 4: Probability of the sabotage action at all states in the sabotage game for ROTATE teammates trained with per-state regret (ours) vs per-trajectory regret. Results are aggregated across three trials.
RQ2b: Does per-state regret lead to improved generalization compared to per-trajectory regret? (Yes)
RQ2a demonstrated that ROTATE with per-state regret (ours) leads to teammate policies that sabotage less often in an illustrative matrix game, compared to ROTATE with per-trajectory regret. Here, we investigate whether the this translates to improved generalization against the unseen evaluation teammates. All configurations other than the teammate’s policy objective are kept identical, including the data used to train the teammate value functions. Fig. 3(b) shows that ROTATE with per-state regret outperforms ROTATE with per-trajectory regret on all tasks except AA, confirming the superiority of per-state regret. As discussed in RQ1, we observe that AA is a layout where an ego agent is less susceptible to sabotage, due to the separated kitchen layout. App. C.2 presents additional experiments testing ROTATE with CoMeDi-style mixed-play rollouts, and alternative methods to compute per-state regret—ultimately finding that ROTATE outperforms all variations.